# Nanchong, Sichuan 637002, People’s Republic of China

### v1

@@ 4 Dynamical evolution
For the convenience of numerical calculations, using the light-cone coordinates $u=t-r_{*}$ and $v=t+r_{*}$, the wave equation (52) can be rewritten as
@@ 5 Conclusions and outlooks
  • Thermodynamical quantities:
    The thermodynamic quantities of the static neutral AdS black hole in Case A in the KR gravity were calculated and obeyed both the first law and the Bekenstein-Smarr mass formula.
  • RII:
    The static neutral AdS black hole of Case A in the KR gravity, which can violate the RII, marks the first discovery of a black hole in static AdS spacetime that can be superentropic due to Lorentz violation; however, Case B always obeys the RII.
  • Dynamical evolution:
    The dynamical evolution waveform of scalar fields in static neutral KR spacetimes decays faster and has a shorter lifespan compared to Schwarzschild spacetime, and Case B is more significantly affected.
There are three promising further topics to be pursued in the future. Although the black hole metrics of Case A in the KR gravity theory have a form similar to those of the static black hole in bumblebee gravity theory Casana et al. (2018); Maluf and Neves (2021), their gravitational perturbation master equations will be entirely different because they originate from different field equations. Based on our previous work on black hole perturbations in bumblebee gravity Liu et al. (2023b), we conclude that the gravitational perturbation isospectrality is broken Liu et al. (2023a). However, for the KR gravity theory, this remains an open question. Furthermore, constructing rotating black hole solutions is an intriguing aspect of the KR gravity theory. When black holes have angular momentum, they exhibit additional properties such as superradiant instability Berti et al. (2009) and deformation of the black hole shadow Liu et al. (2024b); Chen and Jing (2024); Zhong et al. (2021); Guo et al. (2018). Nevertheless, solving the gravitational field equations to obtain an exact rotating solution is very challenging. One can consider the slow rotation approximation Pani et al. (2012, 2013), which greatly reduces the difficulty of the solution while still accurately reflecting the properties of black holes with angular momentum. For more details, please see our separate work Liu et al. (2024c) about the slow rotation extension of the present paper. In addition, exploring the topological classifications Wei et al. (2022); Wu (2023a); Wu and Wu (2023); Wu (2023b, c, d); Zhu et al. (2024) and phase transition criticality Hawking and Page (1983); Kubiznak and Mann (2012); Ahmed et al. (2023); Wu et al. (2024) of the static neutral black holes in KR gravity found in this paper is expected to significantly advance our understanding of the intrinsic nature of gravity.
@@ Acknowledgements.
We thank Professor Shuang-Qing Wu for useful discussions. This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12205243, No. 12375053, No. 12122504, and No. 12035005; the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347; the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028; the innovative research group of Hunan Province under Grant No. 2024JJ1006; and the Hunan Provincial Major Sci-Tech Program under Grant No.2023ZJ1010.
@@ Appendix A The covariant derivative of KR fields
$\nabla_{c}\nabla_{a}\beta^{cd}\beta_{bd}=$ $\Gamma^{c}{}_{ad}\Gamma^{e}{}_{ce}\beta_{bf}\beta^{df}-\Gamma^{c}{}_{ad}\Gamma^{d}{}_{ce}\beta_{bf}\beta^{ef}+\Gamma^{c}{}_{ad}\Gamma^{d}{}_{be}\beta_{cf}\beta^{ef}-\Gamma^{c}{}_{ab}\Gamma^{d}{}_{de}\beta_{cf}\beta^{ef}$ (57)
$+\Gamma^{c}{}_{cd}\beta^{de}\partial_{a}\beta_{be}-\Gamma^{c}{}_{bd}\beta^{de}\partial_{a}\beta_{ce}+\Gamma^{c}{}_{cd}\beta_{be}\partial_{a}\beta^{de}-\Gamma^{c}{}_{bd}\beta_{ce}\partial_{a}\beta^{de}$
$-\beta_{ce}\beta^{cd}\partial_{d}\Gamma^{e}{}_{ab}-\partial_{a}\beta^{cd}\partial_{d}\beta_{bc}-\partial_{a}\beta_{bc}\partial_{d}\beta^{cd}-\beta^{cd}\partial_{d}\partial_{a}\beta_{bc}-\beta_{bc}\partial_{d}\partial_{a}\beta^{cd}$
$-\beta_{bc}\beta^{cd}\partial_{e}\Gamma^{e}{}_{ad}+\Gamma^{c}{}_{ab}\beta^{de}\partial_{e}\beta_{cd}+\Gamma^{c}{}_{ab}\beta_{cd}\partial_{e}\beta^{de},$
$\nabla^{c}\nabla_{c}\beta_{a}^{~{}d}\beta_{bd}=$ $\Gamma^{c}{}_{be}\Gamma^{ed}{}_{d}\beta_{a}{}^{f}\beta_{cf}+\Gamma^{c}{}_{ae}\Gamma^{ed}{}_{d}\beta_{bf}\beta_{c}{}^{f}+\Gamma^{c}{}_{b}{}^{e}\Gamma^{d}{}_{ce}\beta_{a}{}^{f}\beta_{df}+\Gamma^{c}{}_{a}{}^{e}\Gamma^{d}{}_{be}\beta_{c}{}^{f}\beta_{df}$
$+\Gamma^{c}{}_{ae}\Gamma^{d}{}_{b}{}^{e}\beta_{c}{}^{f}\beta_{df}+\Gamma^{c}{}_{a}{}^{e}\Gamma^{d}{}_{ce}\beta_{bf}\beta_{d}{}^{f}-\Gamma^{ce}{}_{e}\beta_{bd}\partial_{c}\beta_{a}{}^{d}-\Gamma^{ce}{}_{e}\beta_{a}{}^{d}\partial_{c}\beta_{bd}$
$+\beta_{bc}\beta^{c}{}_{e}\partial^{d}\Gamma^{e}{}_{ad}+\beta_{a}{}^{c}\beta_{ce}\partial^{d}\Gamma^{e}{}_{bd}-\Gamma^{c}{}_{b}{}^{e}\beta_{cd}\partial_{e}\beta_{a}{}^{d}-\Gamma^{c}{}_{a}{}^{e}\beta_{c}{}^{d}\partial_{e}\beta_{bd}$
$-\Gamma^{c}{}_{b}{}^{e}\beta_{a}{}^{d}\partial_{e}\beta_{cd}-\Gamma^{c}{}_{a}{}^{e}\beta_{bd}\partial_{e}\beta_{c}{}^{d}+\partial_{e}\beta_{bc}\partial^{e}\beta_{a}{}^{c}-\Gamma^{c}{}_{be}\beta_{cd}\partial^{e}\beta_{a}{}^{d}$
$+\partial_{e}\beta_{a}{}^{c}\partial^{e}\beta_{bc}-\Gamma^{c}{}_{ae}\beta_{c}{}^{d}\partial^{e}\beta_{bd}-\Gamma^{c}{}_{be}\beta_{a}{}^{d}\partial^{e}\beta_{cd}$
$-\Gamma^{c}{}_{ae}\beta_{bd}\partial^{e}\beta_{c}{}^{d}+\beta_{bc}\partial^{e}\partial_{e}\beta_{a}{}^{c}+\beta_{a}{}^{c}\partial^{e}\partial_{e}\beta_{bc}.$
@@ Appendix B CFM for the scalar fields
To verify the accuracy of the numerical method in Sec. 4, we will use the CFM to calculate, as reference data, the eigenfrequencies (i.e., QNMs) of the scalar field equation (52), following the Leaver method Leaver (1985, 1986); Pani (2013); Liu et al. (2023c). We need to transform the problem from the time domain to the frequency domain by setting $Z(t,r)=e^{-i\omega t}Z(r)$. For both Case A ($s=0$) and Case B ($s=1$), Eq. (52) can be uniformly rewritten as
@@
$\alpha_{n}$ $=n\left[n-2ir_{h}\sqrt{(1-\ell)^{s+1}}\omega\right],\qquad\gamma_{n}=\left[1+n-2ir_{h}\sqrt{(1-\ell)^{s+1}}\omega\right]^{2},$ (65)
$\beta_{n}$ $=8r_{h}^{2}(1-\ell)^{s+1}\omega^{2}-2n^{2}-L(L+1)(1-\ell)+(1+2n)\left[4ir_{h}\sqrt{(1-\ell)^{s+1}}\omega-1\right].$

### v2

@@ IV Dynamical evolution
For the convenience of numerical calculations, using the light-cone coordinates $u=t-r_{*}$ and $v=t+r_{*}$ PRD109-104060 , the wave equation (55) can be rewritten as
@@ V Conclusions and outlooks
  • Thermodynamical quantities: The thermodynamic quantities of the static neutral AdS black hole in Case A in the KR gravity were calculated and obeyed both the first law and the Bekenstein-Smarr mass formulas.
  • RII: The static neutral AdS black hole of Case A in the KR gravity, which can violate the RII, marks the first discovery of a black hole in static AdS spacetime that can be superentropic due to Lorentz violation; however, Case B always obeys the RII.
  • Dynamical evolution: The dynamical evolution waveform of scalar fields in static neutral KR spacetimes decays faster and has a shorter lifespan compared to Schwarzschild spacetime, and Case B is more significantly affected.
There are three promising further topics to be pursued in the future. Although the black hole metrics of Case A in the KR gravity theory have a form similar to those of the static black hole in bumblebee gravity theory Casana et al. (2018); Maluf and Neves (2021), their gravitational perturbation master equations will be entirely different because they originate from different field equations. Based on our previous work on black hole perturbations in bumblebee gravity Liu et al. (2023b), we conclude that the gravitational perturbation isospectrality is broken Liu et al. (2023a). However, for the KR gravity theory, this remains an open question. Furthermore, constructing rotating black hole solutions is an intriguing aspect of the KR gravity theory. When black holes have angular momentum, they exhibit additional properties such as superradiant instability Berti et al. (2009) and deformation of the black hole shadow Liu et al. (2024b); Chen and Jing (2024); Zhong et al. (2021); Guo et al. (2018). Nevertheless, solving the gravitational field equations to obtain an exact rotating solution is very challenging. One can consider the slow rotation approximation Pani et al. (2012, 2013), which greatly reduces the difficulty of the solution while still accurately reflecting the properties of black holes with angular momentum. For more details, please see our separate work Liu et al. (2024c) about the slow rotation extension of the present paper. In addition, exploring the horizon geometry PRD103-104020 ; PRD104-L121501 , topological classifications Wei et al. (2022); Wu (2023a); Wu and Wu (2023); Wu (2023b, c, d); Zhu et al. (2024), and phase transition criticality Hawking and Page (1983); Kubiznak and Mann (2012); Ahmed et al. (2023); Wu et al. (2024) of the static neutral black holes in KR gravity found in this paper is expected to significantly advance our understanding of the intrinsic nature of gravity.
@@ Acknowledgements.
We are greatly indebted to the anonymous referee for his/her constructive comments to improve the presentation of this work. We also thank Professor Shuang-Qing Wu for useful discussions. This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12205243, No. 12375053, No. 12122504, and No. 12035005; the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347; the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028; the innovative research group of Hunan Province under Grant No. 2024JJ1006; and the Hunan Provincial Major Sci-Tech Program under Grant No.2023ZJ1010.
@@ Appendix A The covariant derivative of KR fields
$\nabla_{c}\nabla_{a}\beta^{cd}\beta_{bd}=\Gamma^{c}{}_{ad}\Gamma^{e}{}_{ce}\beta_{bf}\beta^{df}-\Gamma^{c}{}_{ad}\Gamma^{d}{}_{ce}\beta_{bf}\beta^{ef}$ (60)
$+\Gamma^{c}{}_{ad}\Gamma^{d}{}_{be}\beta_{cf}\beta^{ef}-\Gamma^{c}{}_{ab}\Gamma^{d}{}_{de}\beta_{cf}\beta^{ef}+\Gamma^{c}{}_{cd}\beta^{de}\partial_{a}\beta_{be}$
$-\Gamma^{c}{}_{bd}\beta^{de}\partial_{a}\beta_{ce}+\Gamma^{c}{}_{cd}\beta_{be}\partial_{a}\beta^{de}-\Gamma^{c}{}_{bd}\beta_{ce}\partial_{a}\beta^{de}$
$-\beta_{ce}\beta^{cd}\partial_{d}\Gamma^{e}{}_{ab}-\partial_{a}\beta^{cd}\partial_{d}\beta_{bc}-\partial_{a}\beta_{bc}\partial_{d}\beta^{cd}$
$-\beta^{cd}\partial_{d}\partial_{a}\beta_{bc}-\beta_{bc}\partial_{d}\partial_{a}\beta^{cd}-\beta_{bc}\beta^{cd}\partial_{e}\Gamma^{e}{}_{ad}$
$+\Gamma^{c}{}_{ab}\beta^{de}\partial_{e}\beta_{cd}+\Gamma^{c}{}_{ab}\beta_{cd}\partial_{e}\beta^{de},$
$\nabla^{c}\nabla_{c}\beta_{a}^{~{}d}\beta_{bd}=\Gamma^{c}{}_{be}\Gamma^{ed}{}_{d}\beta_{a}{}^{f}\beta_{cf}+\Gamma^{c}{}_{ae}\Gamma^{ed}{}_{d}\beta_{bf}\beta_{c}{}^{f}$ (61)
$+\Gamma^{c}{}_{b}{}^{e}\Gamma^{d}{}_{ce}\beta_{a}{}^{f}\beta_{df}+\Gamma^{c}{}_{a}{}^{e}\Gamma^{d}{}_{be}\beta_{c}{}^{f}\beta_{df}+\Gamma^{c}{}_{ae}\Gamma^{d}{}_{b}{}^{e}\beta_{c}{}^{f}\beta_{df}$
$+\Gamma^{c}{}_{a}{}^{e}\Gamma^{d}{}_{ce}\beta_{bf}\beta_{d}{}^{f}-\Gamma^{ce}{}_{e}\beta_{bd}\partial_{c}\beta_{a}{}^{d}-\Gamma^{ce}{}_{e}\beta_{a}{}^{d}\partial_{c}\beta_{bd}$
$+\beta_{bc}\beta^{c}{}_{e}\partial^{d}\Gamma^{e}{}_{ad}+\beta_{a}{}^{c}\beta_{ce}\partial^{d}\Gamma^{e}{}_{bd}-\Gamma^{c}{}_{b}{}^{e}\beta_{cd}\partial_{e}\beta_{a}{}^{d}$
$-\Gamma^{c}{}_{a}{}^{e}\beta_{c}{}^{d}\partial_{e}\beta_{bd}-\Gamma^{c}{}_{b}{}^{e}\beta_{a}{}^{d}\partial_{e}\beta_{cd}-\Gamma^{c}{}_{a}{}^{e}\beta_{bd}\partial_{e}\beta_{c}{}^{d}$
$+\partial_{e}\beta_{bc}\partial^{e}\beta_{a}{}^{c}-\Gamma^{c}{}_{be}\beta_{cd}\partial^{e}\beta_{a}{}^{d}+\partial_{e}\beta_{a}{}^{c}\partial^{e}\beta_{bc}$
$-\Gamma^{c}{}_{ae}\beta_{c}{}^{d}\partial^{e}\beta_{bd}-\Gamma^{c}{}_{be}\beta_{a}{}^{d}\partial^{e}\beta_{cd}-\Gamma^{c}{}_{ae}\beta_{bd}\partial^{e}\beta_{c}{}^{d}$
$+\beta_{bc}\partial^{e}\partial_{e}\beta_{a}{}^{c}+\beta_{a}{}^{c}\partial^{e}\partial_{e}\beta_{bc}.$
@@ Appendix B CFM for the scalar fields
To verify the accuracy of the numerical method in Sec. IV, we will use the CFM to calculate, as reference data, the eigenfrequencies (i.e., QNMs) of the scalar field equation (55), following the Leaver method Leaver (1985, 1986); Pani (2013); Liu et al. (2023c). We need to transform the problem from the time domain to the frequency domain by setting $Z(t,r)=e^{-i\omega t}Z(r)$. For both Case A ($s=0$) and Case B ($s=1$), Eq. (55) can be uniformly rewritten as
@@
$\alpha_{n}$ $=n\left[n-2ir_{h}\sqrt{(1-\ell)^{s+1}}\omega\right],$ (69)
$\beta_{n}$ $=8r_{h}^{2}(1-\ell)^{s+1}\omega^{2}-2n^{2}-L(L+1)(1-\ell)$
$\quad+(1+2n)\left[4ir_{h}\sqrt{(1-\ell)^{s+1}}\omega-1\right],$
$\gamma_{n}$ $=\left[1+n-2ir_{h}\sqrt{(1-\ell)^{s+1}}\omega\right]^{2}.$